OpenStudy (anonymous):

A photograph now measures 33 cm wide by 45 cm long. Before it was enlarged, its width was 11 cm. Which equation represents this proportion correctly? sorry the a b c d wont show up

OpenStudy (anonymous):

33:45 BEFORE IT WAS 11:15

OpenStudy (anonymous):

What do you mean the answer is 11:15?

OpenStudy (anonymous):

11/N = 33/45 11 (45/33) = N 45/3 = N 15 = N

OpenStudy (anonymous):

I SUPOSE THE EQUATION IS 11/N = 33/45 =)